40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Nov 19 17:12:22 2010, yuta, Update, CDS, ETMX damped, but with weird TO_COIL matrix ETMXdamping.png
    Reply  Mon Nov 22 12:00:18 2010, josephb, Update, CDS, Updated Computer Restart Procedures for FB 
Message ID: 3957     Entry time: Fri Nov 19 17:12:22 2010     Reply to this: 3962
Author: yuta 
Type: Update 
Category: CDS 
Subject: ETMX damped, but with weird TO_COIL matrix 

Background:
  c1iscex machine is currently being setup and RT model c1scx is running.
  But ETMX(south) didn't seem to be damped, so I checked it.

What I did:
  1. Checked the wiring. It seemed to be OK.
    Looked LEMO monitor output of SUS PD Whitening Board(D000210) with oscilloscope and they seemed to be getting some sensor signal except SDSEN.
      SDSEN is funny. C1:SUS-ETMX_SPDMon decreases slowly when PD input cable is disconnected, and increases slowly when connected.
      There might be some problem in the circuits.
    Looked LEMO monitor output of SOS Coil Driver Module(D010001) with oscilloscope and they seemed to be receiving correct signal from DAC.
      When ULCOIL offset is added, ch1 increased and so on.

  2. Checked the direction of SUSDOF motion when kicked with one coil.
    The result was;

kick (+) POS PIT YAW
ULCOIL + + +
URCOIL + - +
LRCOIL + - -
LLCOIL + + -


    This table tells you, when ULCOIL_OFFSET increases, SUSPOS increases and so on.
    If URCOIL and LLCOIL are swapped, they look correct.
    Also, they have opposite sign to the usual optics(e.g. MCs, BS, PRM).

  3. Changed TO_COIL matrix according to the table above(see Attachment #1). Changed signs of XXCOIL_GAINs.

  4. ETMX damped!

Plan:
  - Check the wiring after SOS Coil Driver Module and circuit around SDSEN
  - Check whitening and dewhitening filters. We connected a binary output cable, but didn't checked them yet.
  - Make a script for step 2
  - Activate new DAQ channels for ETMX (what is the current new fresh up-to-date latest fb restart procedure?)

Attachment 1: ETMXdamping.png  40 kB  | Hide | Hide all
ETMXdamping.png
ELOG V3.1.3-